OpenStudy (anonymous):

Solve for x in the proportion 3x over 3x - 36 = -1 over x

OpenStudy (kropot72):

Cross multiply to get: 3x^2 = -3x + 36 3x^2 + 3x -36 = 0 Now solve the quadratic

OpenStudy (kropot72):

Simplify the quadratic by dividing both sides by 3 \[x ^{2}+x-12=0\] Can you factorise this?

OpenStudy (precal):

(x+4)(x-3)=0 finish solving it